### v1

# Beyond the Main Sequence: Binary Evolution Pathways to Blue Stragglers in the Gaia Era

Abstract
Context. The study of blue straggler stars (BSS) provides insight into the mechanisms of stellar mass exchange during binary stellar evolution and the complex gravitational interactions within dense stellar systems. In combination, they enhance our understanding of the possible life cycles of stars and the evolutionary pathways of star clusters.
Aims. We study the populations of BSSs in 41 globular clusters (GCs) and 42 open clusters (OCs) adopting photometry, proper motion, and parallax from the Gaia Data Release 3 (DR3), confirming their cluster membership. We find a total of 4399 BSSs: all GCs show the presence of BSSs (3965 or of the sample), whereas only 42 out of 129 studied OCs show the presence of BSSs (434 or of the sample). Clusters younger than Myr do not host any BSSs.
Methods. We derive their astrophysical parameters such as effective temperature, surface gravity, and mass based on color-temperature relations, isochrone models and Gaia DR3 spectroscopy (if available). We find values for and ; and an average mass of and for GCs and OCs, respectively. We finally compute the difference of the BSS mass and the main sequence turn-off (MSTO) mass of its respective cluster, normalized by the MSTO mass, for every identified BSS.
Results. Based on this parameter and BSSs ages derived from isochrone models, we find: i) GC BSSs that are most likely to be formed through collisions show a “boost” for ages Gyr, in agreement with ages for core-collapse events in GCs reported in previous studies; ii) a double sequence for GC BSSs that could indicate the hint of a pre/post-merger or close-binary scenario.
Key Words.:
blue stragglers – globular clusters: general – open clusters and associations: general1 Introduction
1.1 Overview and Formation Mechanisms
Blue Stragglers Stars (BSSs) are stars that appear to be younger and more massive than the bulk of other stars in their host stellar population. They are characterized by their location in the color-magnitude diagram (CMD), where they appear hotter and bluer than their main-sequence (MS) neighbours.
BSSs were first identified by Sandage (1953) through his observations of the globular cluster (GC) Messier 3 (M3) and some years later they were reported in the open cluster (OC) NGC 7786 by Burbidge & Sandage (1958). Ever since, their presence has been noticed in distinct structures: as field stars (Preston & Sneden, 2000; Jofré et al., 2016), in OB associations (Mathys, 1987), the Galactic bulge (Clarkson et al., 2011) and in dwarf galaxies (Da Costa, 1984; Mapelli et al., 2009).
BSSs put the theory of stellar evolution in “check”: i.e. the more massive a star is, the faster it evolves. For this reason, these stars should have evolved off the MS branch a long time ago; but we still find them populating this branch as an apparent extension of the MS in a CMD. Three scenarios have been proposed to explain their formation: i) Mass-transfer (MT) in binary systems (McCrea, 1964), where a star is receiving material from its companion, extending its life; ii) Stars formed through collisions/mergers (Hills & Day, 1976; Lombardi et al., 2002); and iii) Hierarchical triple systems: magnetized winds or stellar evolution cause the loss of angular momentum in a multiple system, which leads to collisions (or MT) between the stars that conform the system (Perets & Fabrycky, 2009; Naoz & Fabrycky, 2014). It has been shown that the BSS properties scale with different astrophysical parameters of their parent cluster (Knigge et al., 2009; Leigh et al., 2011; Jadhav & Subramaniam, 2021), which, in principle, could help constrain which of the formation pathways dominates. However, it is important to mention that these formation mechanisms most likely coexist, i.e. they operate simultaneously (Mapelli et al., 2006; Leigh et al., 2011).
Ferraro et al. (2009) found a double sequence of BSSs in the GC M30: a blue and a red one, both roughly aligning with the zero-age MS (ZAMS). The authors argue that the distinction between these two BSS sequences may be due to their origin: collisional BSSs (Col-BSSs), which are deemed more massive, hotter and, therefore, appear bluer in the CMD, whereas MT-BSSs would be less massive, mildly cooler than Col-BSSs and, therefore, appear as a redder BSS sequence. This is not an isolated case, double/multiple sequences of BSSs have been reported for more GCs such as NGC 362 (Dalessandro et al., 2013), NGC 1261 (Simunovic et al., 2014), M 15 (Beccari et al., 2019) and NGC 6256 (Cadelano et al., 2022). The explanation given for all these cases leads to a core-collapse scenario of the host GC in which the central interactions between stars increase, resulting in the formation of BSSs. Therefore, a double/multiple BSS sequence could be a vestige of this event. It also has been found that some BSSs do show the presence of a hot companion (Gosnell et al., 2015; Sindhu et al., 2019), providing supporting evidence for the MT formation scenario in action. Linking these concepts, Dattatrey et al. (2023, hereafter D23) have analyzed the spectral energy distribution (SED) for multiple BSSs in the GC NGC 362. They demonstrate that BSSs that do and do not have a hot companion lie on the red and blue BSS sequence, respectively, in agreement with explanations provided by Dalessandro et al. (2013) for the same GC. More recently, a double BSS sequence has also been detected in the OC Berkeley 17 by Rao et al. (2023a). Given the lack of core collapse in OCs, and, therefore, a considerable attenuation of the Col-BSS formation channel, the authors argue that the presence of this double BSS sequence may be due to a significant difference in BSS rotational velocities, the presence of multiple stellar populations (MSP), or due to extinction effects (for more details see Rao et al., 2023a, and references therein).
1.2 Combining Gaia and HST data
The arrival of the Gaia mission (Gaia Collaboration et al., 2016) with data for almost 1.7 billion objects has opened new opportunities to explore the kinematics of our galaxy and has helped to identify and confirm the cluster membership of stars in GCs (Vasiliev & Baumgardt, 2021) and OCs (Cantat-Gaudin et al., 2018; Dias et al., 2021). However, Gaia faces limitations in overcrowded regions – large errors for its astrometric parameters since it cannot reliably resolve sources – such as GC centres. Nevertheless, BSSs studies in crowded regions are covered by HST data, which can lead to high accuracy measurements in the densest regions (e.g. Baldwin et al., 2016) and, also, since HST features a broader SED coverage compared with Gaia, BSSs can be easily identified through the so-called “UV-route” (Ferraro et al., 1997; Raso et al., 2017; Ferraro et al., 2018, and references therein).
It is particularly interesting to analyze the radial distribution of BSSs, especially in the inner regions of clusters. Since BSSs are more massive compared with other stars in the same cluster, they tend to undergo efficient mass segregation, i.e. “sink” to the inner parts of the cluster (Contreras Ramos et al., 2012). It, is therefore, expected that these peculiar stars can be used to measure the dynamical state of a cluster, i.e., they can be used as “dynamical clocks” in GCs (Ferraro et al., 2012, 2023) and OCs (Rao et al., 2021, 2023b).
Gaia Data Release 3 (Gaia Collaboration et al., 2023, hereafter DR3) for OCs does not suffer from overcrowding, even in the core of OCs. For example, using Gaia Data Release 2 (Gaia Collaboration et al., 2018, hereafter DR2), Rain et al. (2021, hereafter R21) created a catalog for 408 BSSs in different OCs, while Jadhav & Subramaniam (2021, hereafter JS21) have studied the scaling relations of BSSs and their parent OCs. Gaia DR2 has also opened new opportunities to study GCs and their kinematics (e.g., Vasiliev & Baumgardt, 2021). In the past, several HST-based studies have characterized BSSs in GCs (Simunovic & Puzia, 2016). However, the limitation of HST-based studies was that it can only cover the inner regions of clusters, i.e. from the center up to about the half-light radius, . For example, the “HST UV Globular Cluster Survey” (Piotto et al., 2015; Nardiello et al., 2018, also known as “HUGS”) data covers this region and does not include the outskirts of the target GCs. Therefore, combining Gaia DR3 with HST will help us to analyze the radial distribution of BSSs in GCs from the core to outer radii and to determine whether the GCs are dynamically settled or not. In the case of OCs, Gaia DR3 alone can be used to study their dynamical evolution. Hence, a homogeneous catalog of BSSs in GCs and OCs is necessary to understand their formation channels and to examine the dynamical evolution of their host clusters.
In this work, we use Gaia DR3 to create a new homogeneous BSSs catalogue of GCs and OCs, in contrast to previous ones (Rain et al., 2021; Simunovic & Puzia, 2016). This work includes BSSs that were previously not covered due to the radial coverage limitations of prior work. This study also provides various newly derived astrophysical parameters of BSSs from color–temperature relation, isochrone models, and DR3 spectra.
This paper is organized as follows. In Section 2, we describe the selection of cluster members and the BSSs selection. We introduce the different methods applied to obtain different astrophysical parameters for our identified BSS cluster members. In Section 3, we present our results for the total number of identified BSSs and their fraction with respect to their parent cluster MS stars, as well as the distributions for the derived astrophysical BSS parameters. We analyze the results and compare them with previous ones in the literature, and discuss the implications. We finally discuss possible uses and extensions for the identified BSS sample. In Section 4, we summarize the main results of this work.
2 Data Selection
To select the clusters that are studied in this work, we first select GCs and OCs that have a significant amount of members based on previous studies. We select all the GCs from Vasiliev & Baumgardt (2021) catalog with at least 1000 identified member stars and all the OCs from Dias et al. (2021) catalog with at least identified 350 member stars. We find 41 GCs and 129 OCs following the above criteria.
2.1 Stars Cluster Membership
To identify cluster members we apply a technique developed and used by Cordoni et al. (2018, hereafter C18), with a minor change for OCs (see step 2.), which has also been used by Cordoni et al. (2020) for GCs. Our procedure is summarized as follows:
1.
First, we select data from Gaia DR3 based on coordinates provided by Vasiliev & Baumgardt (2021) for GCs and Dias et al. (2021) for OCs. As the data query radius we use the Rscale value –the scale radius of Gaia-detected cluster members–, plus , from Vasiliev & Baumgardt (2021) for GCs. For OCs we select stars inside –the radius containing half of the identified members from Dias et al. (2021)–, with a maximum apparent radius of .
2.
We only keep data with G magnitudes in the range between and , with an error in proper motion (in both components along RA and Dec) lower than and the renormalised unit weight error (ruwe) lower than 1.4.
3.
We select stars inside an ellipse (note: C18 uses a circle) in a vector point diagram (VPD), whose center is based on proper motion coordinates provided by Vasiliev & Baumgardt (2021) for GCs and Dias et al. (2021) for OCs. We slightly vary the minor and major axis of this ellipse, as well as its inclination angle with respect to the y-axis in the VPD, and we keep those parameters that maximize the number of stars within the ellipse.
4.
We compute the parameter , defined by C18, as:
(1) where and are the median of the proper motion along RA and DEC components, respectively, for data within the chosen ellipse. In short, is the distance of every star in the VPD-space to the center of the ellipse.
5.
We plot the G magnitude vs. and split the magnitude component into bins with bin size of mag. In every bin, we compute the median values for mag and the parameter, and the rms () for . We then create points with coordinates G median (vertical axis) and (horizontal axis). We interpolate all the points created in this way and we keep all the values within under this interpolation.
6.
We repeat the same procedure as the previous step, but with the parallax () instead of . For every bin, we create two boundaries with coordinates points based median mag (vertical axis) and median (horizontal axis), interpolate them and keep data between these boundaries.
We iterate this procedure a total of 3 times starting from step 3. For a more detailed description of this procedure, we encourage the reader to view section 2 and figure 1 from C18 and Cordoni et al. (2020). Figure 1 shows the distribution of the member stars in our GC and OC sample. We have also compared our cluster members with the literature (Dias et al., 2021; Vasiliev & Baumgardt, 2021; Hunt & Reffert, 2023; Cantat-Gaudin et al., 2020a) and found a very good match with them. The small discrepancy in the crossmatch is found mostly toward the fainter magnitudes (G18). The details of this crossmatch are described in the Appendix A.
2.2 Selection of Blue Straggler Stars
We create a CMD ridge-line for each cluster using the median color and magnitude in bins in the vs. CMD and select MS, sub-giant branch (SGB) and red giant branch (RGB) stars. We prefer over the magnitude in the CMD since it is expected that BSSs will appear brighter at bluer wavelengths across the Gaia SED coverage (see, e.g., Raso et al., 2017). However, the difference of using or does not change our final sample size as covers approximately –, while covers – (Rodrigo et al., 2012; Rodrigo & Solano, 2020), and given the lack of significant near-IR emission of BSS star the wider coverage of the filter mainly adds noise and not useful flux information for BSSs (see e.g. D23, ).
Next, we use PARSEC isochrones (Bressan et al., 2012) that best fit these ridge lines, adopting the stellar population parameters from the literature: For GCs, distances were taken from Baumgardt & Vasiliev (2021), ages from Baumgardt et al. (2023) and extinctions from Harris (2010). All parameters for OCs were taken from Dias et al. (2021), except for two OCs: for Pismin 3 we use parameters from Bisht et al. (2022) and for Trumpler 23 we use parameters from Cantat-Gaudin et al. (2020b). Starting from these literature values, we also generate isochrones with slight input parameter variations to explore the best-fit isochrone morphologies: in both and in steps of . With the color excess–extinction relation (see Schultz & Wiemer, 1975; Cardelli et al., 1989) converted to Gaia extinction coefficients using the relations and (Marigo et al., 2008; Evans et al., 2018). We consider the PARSEC isochrone that best fits our ridge line for each cluster using a least-square fit.
With the best-fitting isochrones, we divide the data below the MS turn-off (MSTO) point –determined by the infinite slope of the best-fitting isochrone– into 10 bins and record the bin with the largest number of stars. For this bin, we compute its root mean square (rms) along the color axis and call it . We repeat the procedure, but now for SGB stars, and measure the rms along the magnitude axis to obtain . To properly select BSSs, we create a region where we shift the best-fitting isochrone to bluer colors and to brighter magnitudes, and adopt an upper cut in magnitude mag above the MSTO for GCs, and 2.5 mags for OCs (Chen & Han, 2009; Leigh et al., 2011).
Finally, we recreate a zero-age main sequence (ZAMS) keeping the stellar population parameters for the best-fitting isochrone and shifting it to bluer colors, i.e. twice the quantity we have shifted the best-fitting isochrone before. The only 2 differences in the BSS selection between GCs and OCs are in the color selection: i) for GCs we have set a cut in the color up to 0.5 mag to colors bluer than the MSTO to avoid selecting evolved horizontal branch (HB) stars (see, e.g., Culpan et al., 2021). We notice that with this criteria, the selected BSS candidates are found to be located at least five sigma away from the BHB ridge line for all GCs. However, we do not apply this criterion to OCs, as they do have any BHB populations. ii) For OCs, we have applied a cut in color to mag to avoid selecting RGB stars for younger clusters.
Finally, we classify stars within this selection region into 3 types: i) if they are bluer than the MSTO they are classified as BSSs; ii) if they are redder than the MSTO and are in the fainter half of the selection region they are classified as yellow straggler stars (YSSs), i.e., evolved BSSs; and iii) if they are redder than the MSTO and lie in the brighter-half of the selection region they are classified as red straggler stars (RSSs), or evolved YSSs. In this work, we will mainly focus on stars classified as BSSs.
A visual representation of our selection criteria can be appreciated in Figure 2 for two representative clusters, GC NGC 6205 and OC NGC 188, where we display the sample selection results along with their respective VPDs.
We detect 4399 BSSs in our cluster sample. All GCs show the presence of BSSs, with a total of 3965 detected BSSs ( of the sample). On the other hand, only 42 out of 129 selected OCs in our sample show the presence of BSSs, with a total of 434 BSSs ( of the sample). Comparing our BSS members with the literature, we find that more than 90% of all BSSs are also classified as members in previous studies.
3 Results and Discussion
3.1 Absolute Numbers and Fractions of BSSs
Figure 3 shows the absolute number () of detected BSSs and their number ratio relative to MS stars () in each of our sample GCs and OCs as a function of the parent cluster age. For the denominator of , we use the number of MS stars in the luminosity range between the MSTO and MSTO+1 mag, similar to previous studies (JS21, Cordoni et al. 2023). We select MS stars using the same criteria for all clusters shown in the bottom panel of Figure 3 to guarantee a homogeneous comparison. Due to the larger average distances of GCs and, therefore, their fainter MSTOs, our earlier photometric depth cut (see Sect. 2.2) collides with the fainter bound of the luminosity range definition for some distant GCs. For this reason we have inspected by eye every CMD in our GC sample and only provide values for those clusters whose MS stars at fainter than the MSTO are within the photometric depth bounds. With this limitation, we are able to obtain values for 21 of 41 sample GCs, which is the reason why the number of GCs in the bottom panel of Figure 3 is smaller than the corresponding number in the upper panel. We have also listed the values of , and in Table 2 along with other stellar parameters of the clusters.
The inset panels show zoom-ins at the oldest ages to better visualize any localized trends for the GC sub-sample. In Figure 3, we also compare the number of BSSs identified in this work against other Gaia-based studies from the literature, i.e. R21, JS21, Li et al. (2023) and Cordoni et al. (2023). We point out that our sample is the first that homogeneously studies BSSs in GCs and OCs. We observe that in all studies, including ours, the BSS population in OCs tends to increase in absolute and fractional numbers towards older cluster ages. However, this trend appears to reach a maximum at GC ages (), and might even be consistent with the numbers of identified BSSs decreasing with increasing age slightly. We also perform sliding histograms of the distributions of () and (dashed line) in Figure 3. shows a clear trend of increasing with age, while the distribution appeared as a plateau in the range before suddenly rising again with the varying BSS fraction at older ages. We also perform the Pearson’s r-coefficients in the Monte Carlo method 111 Pearson’s test measures whether a pair of two variables are linearly correlated or not. The r-coefficient is defined as the covariance of the variables divided by the product of their standard deviations, and the value ranges from 1 to 1. 1 signifies perfect positive linear correlation between the variables, 1 signifies perfect linear anti-correlation, and 0 means no linear correlation. Based on Cohen (1988), coefficient values 0.5 (0.5) indicate a strong positive (negative) linear relation, values between 0.3 and 0.5 (0.3 and 0.5) indicate a moderate positive (negative) linear relation, and values between 0 and 0.3 (0 and 0.3) indicate a weak positive (negative) linear relation. The Pearson’s test also provides a p-value that indicates whether the observed correlation is statistically significant or not. It depends on the number of data points or the sample size. A smaller p-value (generally 0.05) suggests that the correlation is statistically significant, while for larger values (0.05), we can not conclude a significant correlation (Wasserstein & Lazar, 2016). (to include observational errors) for both distributions (Curran, 2014; Privon et al., 2020)222https://github.com/privong/pymccorrelation?tab=readme-ov-file and find that the values are 0.67 for and 0.45 for .333The numbers in all cases denote the correlation coefficient 50th percentile. The lower and upper errors denote the 16th and 75th percentiles, respectively. The p-values for both cases are found to be 0. The coefficients and p-values suggest that and , both correlate with the age of the clusters. Although the correlation for is moderate, one cannot deny the sudden rise of BSS fraction at . The plateau in in the range could be because, at a young age, massive stars reach larger radii in the AGB phase and mass transfer is possible even in wide binaries, which favours BSS formation to make a constant BSS fraction with respect to the MS population (Leiner & Geller, 2021). Other possible mechanisms of more BSS formation are wind accretion from the AGB companion and stable mass transfer in the subgiant branch or Hertzprung gap in younger clusters. However, the increase in BSS fractions at older ages could be due to the longer MS lifetime of low-mass BSSs in older clusters and/or the contribution from collisional BSSs (Leiner & Geller, 2021, see their Fig.5).
In the inset, we highlight the distribution of for GCs and compare the trend with that of OCs. We notice that for OCs, as they get older, the values increase, whereas for GCs, as they become older, this fraction tends to decrease. We also added a linear fit to the inset plot as a dashed red line that supports this observation. We find the best fit to be:
However, Pearson’s r-coefficient () suggests that the anti-correlation is rather weak. A similar weak anti-correlation () is also found for as well. The p-values are found to be 0 for both cases. As the very central regions of the GCs are excluded in this study because of Gaia’s limitation, incorporating HST observations alongside Gaia data would be beneficial for investigating this correlation. At the same time, we need to include more number of GCs with deep photometry, as we missed to include some GCs due to the magnitude limitation of Gaia.
We note that the dynamic range in covered by OCs and GCs is similar, unlike the absolute number of BSSs, where GCs show larger absolute numbers due to their larger masses, i.e. more numerous stars. More specifically, GCs show while for OCs we find a slightly larger mean ratio of .
3.2 BSS Astrophysical Parameters
We derive the following astrophysical parameters for our sample BSSs: effective temperature (), stellar mass (), luminosity () and surface gravity (). We use three different approaches: the infrared flux method (IRFM; from color-temperature relations), the isochrone best–fit method (from PARSEC isochrones) and the XPP from Gaia DR3 spectra, if available. While the isochrone method provides all astrophysical parameters, the IRFM method provides and the XPP method gives & . We mainly use the parameters obtained from the isochrone method however, it is necessary to compare values, obtained from all three methods to check the consistency in the parameter estimations. In Appendix B, we describe all three methods and the comparison of the values obtained from them, in detail. In the following subsections, we summarize the results and compare the stellar parameters with the literature values.
3.2.1 Effective Temperature
We notice that values extracted from different approaches range from 6000 K to 9000 K and appear consistent with each other with a little systematic offset towards higher temperature ( ). From these comparisons, we estimate that the systematic uncertainty between various methods is of the order K. From the IRFM we find for GCs and for OCs. From isochrone fitting we find for GCs and for OCs. Only 537 out of 4399 BSSs ( of our sample) have XPP parameters measured for 279 stars in GCs and 258 stars in OCs. With this method, we find the mean BSS effective temperature of for GCs and for OCs. We have listed these parameters in Table 3 and the full table will be available online in an electronic format.
Table 1 compare our derived values for with those found in the literature for various OCs and GC, as well as BSS samples. We find that our effective temperatures derived are in very good agreement with previous studies. For example, Jadhav et al. (2019), Sindhu et al. (2019) and Pandey et al. (2021) have analyzed different BSSs in the OC NGC 2682 (also known as M67) by fitting models to spectral energy distributions (SEDs) searching for BSS hot companions; these studies derive BSS temperatures in the range between and , which are in excellent agreement with our measurements, yielding a mean for all methods combined. We also point out the excellent agreement with the values derived in Gosnell et al. (2015), where the authors find for 19 BSSs, which is in very good agreement with our results independent of the applied method ( and , for the IRFM and isochrone method, respectively). Finally, our is also in very good agreement with theoretical predictions for BSSs in GCs from Stȩpień & Kiraga (2015), who calculate , and who had applied different initial conditions to binary systems.
3.2.2 Stellar Mass and Age of the BSSs
Stellar mass and age are important parameters in the study of the evolutionary state of BSSs. We derive these parameters along with equivalent stellar luminosity, and stellar surface gravity for all BSSs in our sample based on the isochrone fitting method. We obtain the mean luminosity for GCs and for OCs, and find the mean surface gravity of for GCs and for OCs. These values correspond to a mean stellar mass for BSSs in GCs and for BSSs in OCs, which altogether translates to a mean stellar age of for BSSs in GCs and for BSSs in OCs.
We compare these values with previous results from the literature in Table 1. We find that the average mass of our BSS sample in GCs () differs from previous works, which report (Shara et al., 1997) and (Baldwin et al., 2016). This may be partly explained by the lower BSS numbers of previous studies, but also by either of the previous arguments, i.e. i) limitations of the Gaia data selection and ii) BSS mass segregation in GCs. In particular, the latter argument lends credence to the reason for this offset as the two previous studies are based on HST observations which target the GC centres. Our sample, on the other hand, is based on BSS samples from outside the half-light radius. Nevertheless, as can be appreciated in Table 1, our BSS mass results are in very good agreement with those measured by Simunovic (2017, ), Fiorentino et al. (2014, ) and Stȩpień & Kiraga (2015, ). The comparison with Simunovic (2017) is in particular striking, as this work measured the parameters for 1507 BSSs in 38 GCs using the isochrone fitting technique similar to our procedure, which makes it statistically the most robust comparison.
| Name | Reference | |||||
| Open Clusters | ||||||
| NGC 188 | 19 | — | — | Gosnell et al. (2015) | ||
| 21 | — | — | This Work – IRFM | |||
| 21 | This Work – isochrone | |||||
| NGC 2682 | 21 | — | — | Jadhav et al. (2019) | ||
| 3 | — | — | Sindhu et al. (2019) | |||
| 6 | — | Pandey et al. (2021) | ||||
| 9 | — | — | This Work – IRFM | |||
| 9 | This Work – isochrone | |||||
| NGC 7789 | 15 | — | Vaidya et al. (2022) | |||
| 16 | — | — | This Work – IRFM | |||
| 16 | This Work – isochrone | |||||
| Multiple OCs | 1368 | — | — | Jadhav & Subramaniam (2021) | ||
| 434 | — | — | This Work – IRFM | |||
| 434 | This Work – isochrone | |||||
| Globular Clusters | ||||||
| NGC 362 | 26 | — | Dattatrey et al. (2023) | |||
| 137 | — | — | This Work – IRFM | |||
| 137 | This Work – isochrone | |||||
| NGC 5272 | 7 | — | De Marco et al. (2005) | |||
| 210 | — | — | This Work – IRFM | |||
| 210 | This Work – isochrone | |||||
| NGC 6541 | 9 | — | — | Fiorentino et al. (2014) | ||
| 52 | This Work – isochrone | |||||
| Multiple GCs | 64 | — | Stȩpień & Kiraga (2015) | |||
| 598 | — | — | Baldwin et al. (2016) | |||
| 1507 | — | — | Simunovic (2017) | |||
| 3965 | — | — | This Work – IRFM | |||
| 3965 | This Work – isochrone |
Notes. For data provided by Jadhav & Subramaniam (2021, or JS21) we consider stars labeled as “‘BSS” and “‘pBSS” (probable BSS).
The average BSS mass for our OC sample ( compares favourably with the corresponding value found by JS21 considering the uncertainties. They report an average-mass value of if only confirmed BSSs from their catalog are considered, and if both confirmed and candidate/probable BSSs listed in their catalog are included. As was explained previously, the difference between these works is likely the result of our Gaia magnitude cut-off criterion, discarding all bright stars with as they are likely saturated and, therefore, their Gaia parameters likely affected by systematic errors. This not only explains the difference between our results, as well as why our dispersion in BSS mass for OCs is smaller, but also the lack of BSSs younger than Myr in our sample, most of which would naturally be among the most massive BSSs.
3.3 Effective temperature vs cluster age
In Figure 4 we compare the average MSTO offset in of BSS in a cluster against the corresponding cluster age. To compare self-consistently all the BSSs with respect to the same evolutionary reference point, given the clusters’ very different ages and metallicities, we use the of their MSTO as reference. For each BSS, we obtain . We then compute which is the mean difference between the BSS temperatures and the in each cluster. As XPP parameters are not available for stars at the MSTO, we only do this for obtained by isochrones and color–temperature relation, i.e. IRFM, and recall the systematic offsets found in the comparison of these methods (see Fig. 14). The errorbars in represent the dispersion of BSS temperatures in each cluster, while the errorbars on the horizontal axis are age uncertainties.
In the top panel of figure 4, BSSs in OCs show on average a trend against the parent cluster age where younger OCs show larger offsets, i.e. hotter BSSs, with respect to those found in older OCs, or in other words, the difference between the effective temperature of BSSs and the MSTO effective temperature decreases with increasing parent cluster age. Pearson’s coefficients for IRFM (isochrone) method is (), suggesting the correlation to be weak, which could be due to a large dispersion in the BSS temperatures in the younger cluster. However, we expect the observed trend since younger clusters host more massive stars which are also larger in diameter (e.g. Iben, 2013), and therefore, have likely more efficient MT between binary members and higher collision/capture probabilities with other stars. These hotter BSS stars evolve faster than BSS stars around the MSTO and, thus, the difference in average between BSSs and BSSs around MSTO is expected to decrease with time. We also notice that the dispersion in decreases with the age of the OC, which may indicate that BSSs tend to be more concentrated in the color-luminosity space as the OC ages, which is also shown in the bottom panel of the figure. The Pearson’s coefficient values further confirm the trend: and for IRFM and isochrone methods, respectively.
We do not observe any of these correlations for BSSs in GCs, which is also supported by the Pearson’s coefficient for IRFM (isochrone) method () (see inset panels in Fig. 4). However, we must keep in mind that we might be exposed to biases for the GC sub-sample: i) we have applied a color cut when selecting BSSs in GCs to avoid including BHB stars (see Fig. 2 and Sect. 2.2) which, although unlikely may cut into the distribution of BSSs with extremely large (this could in principle be tested with UV photometry); ii) Gaia is only capable to observe stars outside (half-light radius) in GCs, which could change the mean if we assume that mass segregation has already acted on the BSS population (Contreras Ramos et al., 2012; Ferraro et al., 2012). Without consideration of these potential biases and taking the distributions in Figure 4 at face value, this means that as GCs age past Gyr, BSSs are more or less equally distributed in .
In the middle panel of figure 4, we check if fractional excess with respect to has any correlation with the cluster age. We do not find any trend for either OCs or GCs, which is also supported by the Pearson’s coefficient test. Pearson’s coefficients for IRFM (isochrone) are () for OCs and ().
3.4 BSS Fractional Mass Excess
BSSs are expected to gain mass during their MS lifetime. This means that the “excess” between their current mass and the stellar mass at the MSTO could provide hints about their formation pathways. Recall that, based on previous observations (Dalessandro et al., 2013; Dattatrey et al., 2023), hotter (bluer), but relatively faint BSSs are likely to be formed through collisions, whereas slightly cooler (redder), but slightly brighter BSSs are likely to be formed by MT. Even though this is almost certainly not strictly followed by the universe (modulo stellar-mass loss rate, binary parameters, environmental factors, such as tidal shocks and disruption of binaries, etc.), it could be considered a first scenario for the formation of particular BSS sub-populations. Trying to provide a quantity that could give a hint about which pathway dominate certain evolutionary phases, JS21 defined a parameter called the “fractional mass excess”, or . This quantity is defined as
where is the mass of the BSS and is the stellar mass at the MSTO, which in our case is obtained from isochrone models. In simple words, is the excess between the BSS mass and MSTO mass, normalized by the MSTO mass. Note that is equivalent to , and corresponds to . By definition, this “fractional mass excess” is roughly equivalent to the MT efficiency if we assume that both member stars of the system, the donor and the secondary star, are MS stars (Shao & Li, 2016). JS21 decided to divide the range into three ad-hoc regimes: low– (), high– () and extreme– () where, as a first approximation they speculated, different BSS formation pathways would dominate, which are MT, stellar collisions, and multiple-mergers/MT, respectively.
To make headway in the discussion of BSS formation pathways, we study the behavior of as a function of cluster age and the equivalent BSS age from isochrone fitting, both of which are displayed in Figure 5 along with the BSS mass and mass offset scalings. We focus first on the left-side panels. In the top panel, we observe that, as expected, younger clusters host BSSs with higher masses, which decrease with cluster age. The Pearson coefficient also provide a strong correlation value of , combining both OCs and GCs. This correlation is consistent with the decreasing MSTO mass as a stellar population ages, since the lifetime of a star, , scales as , i.e. more massive stars have shorter lifetimes compared to less massive stars, which translates into the BSS mass relation in the top left panel of Figure 5. This correlation also appears stronger when only OCs are considered (), but this is not true for GC sample (). The reasons for GCs are the same as discussed in the previous sections.
The middle panel on the left side of Figure 5 shows that the difference between the BSS mass and the MSTO mass as a function of cluster age. We see that the BSS with higher mass difference are absent towards the older cluster age with Pearson coefficient of . This is also expected from regular stellar evolution since more massive stars end their life disproportionately earlier while the MSTO mass decreases more slowly with increasing age, which is exactly what is observed, i.e. shrinks as time advances. We also notice here that GC sample alone does not show any correlation ().
More surprisingly, we find no trend of with cluster-age (bottom left panel in Fig. 5) either for OCs or GCs with Pearson coefficient values and , respectively. However, we observe that the fraction of high- () BSSs is larger in OCs () compared to GCs (), which may indicate that the collisional formation channel is more pronounced in younger star clusters. This is turn might be related to sufficiently old dynamical ages of star cluster, which can, in principle, have undergone one or more core collapse, boosting the stellar encounter rates and, therefore, the formation of collision-induced BSS formation. However, this might be accomplished for star clusters at a few Gyr of age, and requires further investigation and modeling. We find and as the BSS fractions with low and high– (split at ), respectively, for OCs. In GCs, we find these proportions to be and for low and high–, respectively. Considering all clusters, the percentage of low– is and for high– we obtain .
Additionally, we plot the BSS mass parameters vs. BSS ages in the right-hand side panels of Figure 5. The relations for and are consistent with those found against the cluster age (left side of the same Figure), i.e. both of these parameters decrease as a function of BSS age and stellar population age. In addition, we see various sequences that indicate that the BSSs in OCs and GCs predominantly populate different parameter spaces. A striking distribution is found in the against plot (bottom right panel) which shows two dominant sequences, both of which are consistent with decreasing as BSSs become older, albeit at different rates. We will later show that these sequences are likely representative of the different BSS formation channels. The pearson coefficients for all these three relations for OCs (GCs) are: () for vs BSS ages; () for vs BSS ages; and () for vs BSS ages.
Figure 6 shows the percentage of low– and high– BSSs as a function of BSS age. For OCs, it is clear that the high– fraction vanishes above BSS ages of , whereas the low– fraction dominates, with a percentage in the same older age range. Curiously, the high– fraction in GCs increases up to for BSS ages between –, while the low– fraction experiences a corresponding dip. This may be due to the predominance of BSS formed through the collisional channel at those ages. Ferraro et al. (2009); Dalessandro et al. (2013); Simunovic et al. (2014) fitted collisional isochrones (Sills et al., 2009) to the blue sequence of BSSs, finding that their ages are , concluding that double sequences of BSSs in GCs may be products of violent interaction episodes in the past core-collapse events, which is in agreement with our age range found for BSS in GCs where the high– fraction increases and the low– fraction dips. For GCs older than the high– fraction rapidly vanishes similar to those for OCs. Overall, we observe for all BSSs in our sample low/high– total fraction is dominated by OCs for younger BSSs ages () and by GCs for older BSSs ages () as expected by their sample characteristics (see Sect. 3.1).
3.5 Comparison with JS21
In Figure 7, we compare our measured values against those from JS21. In the upper panel, the dashed red-line shows the 1-to-1 relation, whereas the yellow solid line marks the best linear fit, with the shaded area representing the confidence interval. Overall, we point out that our values are slightly larger compared with those provided by JS21, as can be appreciated in the inferior panel of Figure 7. An explanation is our Gaia DR3 selection criteria, due to the magnitude selection. We skipped too bright/saturated stars to avoid systematics, which may have been classified as more massive (very bright) BSSs by our models, especially in OCs. In addition, the systematics slope and scatter in Fig. 7 are in part due to JS21 measuring BSS masses using a different isochrone approximation from ours, i.e. for every BSS they estimate its magnitude position on the ZAMS and compute its mass based on this model, not on the best-fit isochrone like in our approach.
We recall that JS21 classified BSSs into three classes, i.e. low– (), high– () and extreme– () objects, which are speculated to predominantly form via MT, stellar collisions, and multiple-mergers/MT, respectively. Apart from our direct comparison of OCs and GCs, the difference between our results and those provided by JS21 is that we were not able to find any stars with in the extreme– regime, most likely due to our Gaia photometry cuts. It is important to mention, and as JS21 have warned, that the “fractional mass excess” does not strictly indicate that how BSSs are formed, but it rather suggests how BSSs may likely be formed. For a deeper analysis, which escapes the Gaia limitations, we do need FUV observations to be able to distinguish the presence or absence of hot companions (e.g. Subramaniam et al., 2016; Sindhu et al., 2019).
We plot the distribution of low- and high– BSSs in OCs and GCs in the Gaia CMD shown in Figure 8, similar to JS21 (see their Fig. 6). We find that BSSs in OCs and GCs classified as high– show a modest difference in color, being bluer than low– BSSs by . The more significant difference is in the absolute magnitude where high– BSSs are, on average, approximately 10 times brighter than low– BSSs. This difference is appreciable for both type of clusters, GCs and OCs. Although this empirical classification resembles the instantaneous mass excess with respect to the MSTO, it does not resemble the classic evolutionary pathways of stars, which leads them from the MS to the colder sub-giant branch and then up the RGB.
3.6 Towards BSS Evolutionary Pathways
With goal of advancing our understanding of the BSS formation channels and capturing the predominance of BSS formation channels, we combine the fractional mass excess parameter () with the our earlier derived equivalent BSS ages. We focus our attention on Figure 9, the upper panel of which reproduces the bottom right panel of Figure 5; however, this time we separate the two sequences444Do not confuse this with the BSS double sequences in a CMD discussed earlier, as the upper panel Figure 9 is not a CMD. into a high-mass excess sequence and low-mass excess through the next procedure:
1.
For the identified BSSs we separate them using as criteria if they belong to an OC or a GC and make a plot vs. (similar as the bottom-right panel of Figure 5).
2.
We then set two edges by eye that enclose all the points in this plane. We define an inferior and superior edge, which is different for OCs and GCs. For OCs we define the lower edge defined as:
(4) where is the BSS age found with isochrone fittings. We then define the superior edge for OCs as:
(5) In a similar way, for GCs we compute these edges as:
(6) and:
(7) -
3.
To check the distribution along the horizontal axis, for every point from OCs in this plane with coordinates , we compute the distance between the point and its superior edge along the vertical axis, and normalize this value by the distance between the superior and inferior edges to “flatten” it. We can reach this using the expression:
(8) -
4.
We repeat the last step, but this time for BSSs from GCs using the respective edges values from step 2.
- 5.
6.
Based on the values that optimize the bimodal distribution for the discovered distances, we compute the probability density function (PDF) of each Gaussian and estimate the membership probability of each point to be a member of each Gaussian.
7.
If the point has a higher probability to belong to the Gaussian with lower distances, this means that the point belongs to the distribution that is closer to the superior edge and, therefore, is classified as a member of the “upper” sequence. Else, they are classified as a member of the “lower” sequence.
To properly compare BSSs with this new classification scheme, we plot them in a Gaia photometry CMD in the lower panel of Figure 9 (bottom panel), which can be directly compared to Figure 8 which encodes the classification from JS21. Our new classification now resembles more the expected stellar evolutionary pathways of BSSs, clearly separating the BSS population in GCs into two dominant sequences. We also observe that the upper- sequence BSSs in OCs populate predominantly the brighter sequence of OC BSSs in the CMD. In general, we find that the color range covered by the two sequences is similar for GCs and OCs, but there is a systematic offset in luminosity with the upper- sequence BSSs appearing about brighter than their lower- sequence counterparts.
In the following, we briefly explore possible explanations and consideration when interpreting these clearly distinct BSS sequences.
3.6.1 Mass Transfer Evolution
From a binary evolution perspective, the upper- sequence may represent a BSS population which was formed through more efficient MT processes. This would imply more accreted mass and, hence, more fuel on the receiving star, accelerating its evolution (McCrea, 1964). Assuming this scenario and a subsequent common-envelope evolution, the separation between the stars of such a binary system will eventually decrease until the final stages which could play out in two scenarios: i) the cores could merge or ii) one of the cores could be ejected from the system through interactions with envelope and/or other stars. The former scenarios is more likely (Hurley et al., 2002). The result of a stellar merger is a massive BSS, which would experience fast subsequent stellar evolution () and be more likely encountered at older BSS ages. We, therefore, interpret the two sequences in the vs. plane as post-merger/close-binary interaction sequence (“upper” sequence) and pre-merger/close-binary interaction sequence (“lower” sequence, see Fig. 9).
3.6.2 Unresolved Binary Stars
It is trivial that some BSSs may be unresolved binary systems (Gosnell et al., 2015; Subramaniam et al., 2016; Sindhu et al., 2019). However, when fitting isochrone models to our BSS sample, we are implicitly assuming single-star evolution. Jiang (2022) has shown through multiple simulations that the total mass of the system is higher in unresolved MT binaries compared with post-MT BSSs by about –. This means that the isochrone fitting method could be overestimating ages and underestimating the masses for unresolved MT BSSs, located preferentially in CMD regions with similar colors as the MSTO and the sub-giant branch. Therefore, the upper- sequence could be stars with an overestimated age –showing higher values– since they are unresolved MT binaries, while the lower- BSSs could be post-MT binaries whose companions have either negligible mass or which are in the post-MT phase.
3.6.3 The Impact of Stellar Radius Variations
As rejuvenated stars evolve of the ZAMS towards the sub-giant branch and are therefore expected to grow in size, we compute the radii of our sample BSSs, , using the relation , where is the gravitational constant, is the stellar mass and is the surface gravity. We assume in cgs units (from isochrone models) and stellar mass in units, computing the value of in solar radii using Astropy routines (Astropy Collaboration et al., 2013). We show the results in Figure 10: and as function of stellar radius (). We note that after adding stellar radius, the separation between upper- and lower- sequences is even more pronounced for GCs.
This is consistent with the expectation of an evolved binary system for BSSs in the upper- sequence (orange squares in Fig. 10), where the system might be in the common-envelope phase. Such systems evolve rapidly and reach subsequent stages with similar characteristics of an SGB or RGB star, where they begin the radial expansion, which qualitatively agrees with our observations. However, we are unable to identify any clear difference between the upper- and lower- sequences for BSSs in OCs, which may be an indication of a different mix of BSS formation mechanisms acting in lower mass, lower density star clusters. We defer the detailed modeling of these systematic differences between OCs and GCs to future work.
4 Conclusions
With the new - classification scheme of BSSs, we inspect the CMDs of all our sample GCs and OCs individually to see whether the BSS sub-populations are differentiated by our classification in the color-magnitude space. We observe that BSSs classified as upper- members are brighter whereas lower- sequence member stars appear fainter. In Figure 11 we present the distribution of upper- and lower- sequence BSSs in the representative GC NGC 362 and OC Berkeley 39. We also overlaid isochrones correspond to the ZAMS and binary-ZAMS. The figure shows that binary-ZAMS passes through majority of the BSSs intersecting both upper- and lower- sequence, while the ZAMS passes through a smaller fraction of lower- sequence (red points). The lack of BSSs on ZAMS also suggests that the lack of collisional BSSs, which is also expected since NGC 362 is a core-collapsed GC and Gaia data are unable to resolve the core of the GCs. We also do not expect to observe collisional BSSs in OCs as the environment is not dense enough to favor collision. Our classification is not similar to the literature classification of blue and red BSSs sequences observed in core-collapsed GCs, where blue BSSs are thought to form through direct collisions, and the BSSs in the red sequence form via MT (Ferraro et al., 2009; Dalessandro et al., 2013). We notice that blue (or collisional) BSSs are dominated by our lower sequence, whereas the red sequence is a mix of upper and lower sequence. Our new BSSs classification based on - mainly highlights the pre- and post- merger/MT scenarios of the BSSs. Our study suggests that the presence of companions across our - classification may imply that most of our stars classified as upper- sequence BSSs could be consistent with the presence of a companion, whereas lower- BSSs may preferentially indicate post-MT/merger products BSSs. This may support our hypothesis that assumes that we could be overestimating BSS ages due to our inability to resolve the binary system and therefore having the BSS classified as upper- sequence member.
The distribution of upper- and lower- sequence BSSs in the CMDs for GCs and OCs is similar for all clusters. The CMDs of all the clusters highlighted with these two sequences are made available online for reference. In some of the clusters (NGC 104, NGC288, NGC 2808, NGC 3201, NGC 5139, NGC 6121), not classified as core-collapsed GCs, we observe a significant number of BSSs are located on the ZAMS or bluer than ZAMS, which most probably form through direct collision or post-MT/merger, as suggested by the literature on core-collapsed GCs. The presence of a significant number of BSSs outside the core or half-light radius of the clusters makes these GCs as interesting candidates and required further detailed analysis to interpret the observed nature.
Much more theoretical work and observations are required to elucidate whether and to what extent our new classification scheme of BSS resembles the formation and evolution mechanisms of these unique stars. For now, we summarize in the following the results and conclusions of our work.
1.
We have selected OCs from Dias et al. (2021) that have more than 350 star members and GCs from Vasiliev & Baumgardt (2021) with more than 1000 star members to ensure well-sampled cluster stellar populations, requesting data from Gaia DR3, for which we adopt an algorithm based on prescriptions from C18 and Cordoni et al. (2020) to extract cluster members.
2.
Using the confirmed cluster members, we fit PARSEC isochrones (Bressan et al., 2012) to obtain parameters such as the age for every cluster in our sample.
3.
Using the ridge line of the observed CMD and the best-fit isochrone found in the previous step along with the ZAMS and the dispersion of stars along these sequences we define a BSS selection. We are able to find BSSs in all our sample GCs and in 42 out of 129 () OCs of our sample. We identify a total of 4399 BSSs; 434 () located in OCs and 3965 () located in GCs.
4.
We extract astrophysical parameters for all identified BSSs using 3 different methods: 1) color–temperature relations (i.e. the infrared-flux method, IRFM); 2) model isochrone fitting; 3) parameters from Gaia DR3 spectra from Zhang et al. (2023, called XPP), if available. From color–temperature relations we only obtain . From isochrone fitting we obtain , , , stellar masses and ages. From XPP we obtain and .
5.
We do not find BSSs in clusters younger than .
6.
We find effective temperatures for BSSs from color–temperature relations and from isochrone–fitting for all our BSS sample. found in OC BSSs are and from color–temperature and isochrone–fitting methods, respectively. BSSs in GCs present slightly lower temperatures with derived from color–temperature relations and from isochrone models. Overall, the difference in found from isochrone–fittings are hotter than those obtained from color–temperature relations. These values are in agreement with previous studies.
7.
Based on the isochrone–fitting method, we find the average mass of all our BSSs to be, approximately, . In OCs, we find the average mass of BSSs at . For GCs we obtain an average BSS mass of . Our average mass for BSSs in OCs are lower to those found by JS21, which is likely to our more conservative selection, however, they are still within their respective error ranges. Our results are in excellent agreement with most other studies.
8.
We compare against its counterpart measured at the MSTO of the parent cluster. Independent of the method applied, for OCs and GCs we find that the difference between the mean BSS and the MSTO decreases with the cluster age. This is in line with expectations from stellar evolution because where massive stars evolve faster than stars at the MSTO. In OCs the dispersion decreases with the cluster age. We do not find this last relation in GCs, which means that the dispersion of BSSs does not depend of the cluster age in GCs.
9.
We follow the definition provided by JS21 for the “fractional mass excess”, or . This parameter quantifies the difference between the BSS mass and the stellar mass at the MSTO –this last based on isochrone-fittings–, normalized by the MSTO mass. is roughly equal/proportional to the MT efficiency in a binary system. These authors classify in three regimes, each suspected to be dominated by a specific BSS formation pathway: i) low– (, likely to be formed via MT), high– (, likely to be formed from mergers/collisions) and extreme– (, likely to be multiple-MT systems). In our sample, we do not find any extreme–, which is likely to be due to the conservative selection, where we avoided stars brighter than –to avoid Gaia systematic errors– that may contain some more massive BSSs. In our final BSS sample, is classified as high– and as low–. For OCs this proportion is and for low– and high–, respectively. For BSSs in GCs, we find these percentages to be for low– and for high–. This suggests that mass-transfer is the most likely formation channel for BSSs; but OCs show a higher fraction of high– BSSs, some of which may have formed through mergers as well.
10.
Finally, we explore a new BSS classification scheme, which uses the vs. . This parameter space shows a clear double sequence for BSS in GCs: one corresponding to an “upper” sequence, with higher as a function of the BSS age, and another “lower” sequence, linked to smaller as a function of the BSS age. We plot BSSs labeled by this classification scheme in the Gaia CMD and find a clear division between these branches; especially for GCs. BSSs on the upper- sequence are brighter than those located on the lower- sequence. We explore two possible explanations: 1) In binary star evolution, mass transfer increases the mass of the receiving star and in the process shrinks the distance between the binaries, while the receiving star begins an accelerated evolution and, hence, will appear as an “older” star in the isochrone tracks. In this scheme, these upper- and lower- sequences are consistent with post-merger/close-binary interaction and pre-merger binary evolution. 2) We might be overestimating the age and underestimating the mass of the upper- sequence BSSs, since we are assuming single star evolution during isochrone fitting. Therefore, the upper- and lower- sequences may be equivalent to an “overestimated” and “realistic” BSSs age, which could be correlated with the presence (or absence) of a BSS companion. Further observations targeting various parameters, such as the stellar rotation speed statistics, chemical compositions and binarity are rquired to shed light on this intriguing new bifurcation in the vs. classification.
All BSS parameters are available from the authors upon reasonable request.
Acknowledgements.
P.K.N. and T.H.P. acknowledge support by the ANID BASAL projects FB210003. T.H.P. gratefully acknowledges support through FONDECYT Regular Project No. 1201016.cluster cluster RA Dec log(age/yr) Av distance [Fe/H] Teff,IRFM Teff,isochrone NBSS NMS fBSS/MS name type (deg) (deg) (mag) (pc) (dex) (K) (K) (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) NGC104 GC 6.175 72.024 10.0700.011 0.124 4367 0.65 590392 5995 415 13691 0.060.002 NGC288 GC 13.381 26.529 10.0640.002 0.031 9000 0.8 6066116 6173 150 – – NGC362 GC 15.885 70.781 10.0680.013 0.155 9090 1.3 6248160 6368 137 1578 0.1710.011 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ Berkeley39 OC 116.893 4.735 9.811 0.029 0.394 4156 0.0 5932 84 6038 22 264 0.14 0.025 Berkeley32 OC 104.738 6.379 9.686 0.022 0.553 3321 0.3 6286 125 6403 14 139 0.151 0.035 Collinder261 OC 189.802 68.492 9.898 0.065 0.999 2706 0.1 5772 72 5865 64 460 0.228 0.025 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ 555The list of columns represented here is: (1) Name of the cluster; (2) whether it is a globular cluster (GC) or open cluster (OC); (3) and (4) represents central coordinates of the cluster, (5) age of the cluster in log scale, (6) extinction (AV) towards the cluster, (7) distance to the cluster, (8) metallicity of the cluster, (9) mean effective temperature of BSS populations in that cluster using IRFM method, (10) mean effective temperature of BSS populations in that cluster using isochrone method; (11) number of BSSs detected in that cluster; (12) number of main sequence (MS) stars in the luminosity range between the MS turn-off (MSTO) and MSTO+1 mag in that cluster; (13) fractional BSSs population to MS populations.
Isochrone Method cluster type Gaia source id Teff,IRFM Teff log(age/yr) M log(L/L⊙) logg Teff,xpp loggxpp Me Mesequence (K) (K) (M⊙) (K) (mag) (mag) (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) (14) (15) NGC2141 OC 3341815582307132416 636382 651059 9.2220.015 1.700.01 1.0400.006 3.8330.010 – – – – 0.17 lower 2.237 0.597 NGC2141 OC 3341815582307133056 639184 646479 9.2220.015 1.720.00 1.0700.003 3.8000.020 – – – – 0.19 lower 2.176 0.589 NGC2141 OC 3341815681086838272 6736101 685595 9.0870.077 1.560.01 0.8000.005 4.1220.025 – – – – 0.08 lower 2.807 0.500 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ NGC104 GC 4689572883825672192 6422126 656746 9.7400.027 1.010.01 0.5500.004 4.1190.015 6831136 4.1420.123 0.28 upper 3.538 0.566 NGC104 GC 4689573433581464192 6262109 638426 9.8420.024 0.930.01 0.3500.003 4.2260.009 6896166 4.2160.2 0.18 lower 4.043 0.609 NGC104 GC 4689574086416430464 6669154 683835 9.5050.035 1.080.01 0.4900.002 4.2710.009 6652159 4.1950.142 0.37 upper 3.651 0.505 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ 666The list of columns presented here is (1) the name of the cluster to which the BSS belong; (2) whether the cluster is classified as an open cluster (OC) or globular cluster (GC); (3) source id of the BSS as per the Gaia DR3; (4) effective temperature of the BSS based on IRFM method; (5) – (9) effective temperature, age, mass, luminosity and surface gravity of the BSS, respectively, estimated using isochrone method; (10) – (11) effective temperature and of the BSS estimated from Gaia-XPP spectrum. (12) fractional mass excess of the BSS based compared to MSTO mass; (13) location of the BSS based on mass excess, as mentioned in Figure 9; (14)-(15) BP magnitude and (BP-RP) colour in absolute scale.
Appendix A Validation of our membership criteria:
To validate our BSS membership criteria as well as all other cluster members, we compare our cluster members with the literature. We only compare the clusters which show the presence of BSSs as we only focus on these clusters in our further analyses. The most recent and extensive membership catalogs using Gaia data are given by Vasiliev & Baumgardt (2021); Hunt & Reffert (2023) for GCs and Cantat-Gaudin et al. (2020a); Hunt & Reffert (2023) for OCs. Vasiliev & Baumgardt (2021) provided membership probability of all stars in 170 GCs using the Gaussian mixture model on Gaia-EDR3 data, where stars with a membership probability greater than 90% are considered to be cluster members. We find more than 90% match in membership for all 41 GCs, with a median crossmatch accuracy of 99.7%. All BSSs detected in our GCs sample are also classified as cluster members in Vasiliev & Baumgardt (2021). Hunt & Reffert (2023) used HDBSCAN algorithm on Gaia DR3 data down to magnitude G 20 mags to recover 7167 clusters, including 134 globular clusters. For GCs, we find that the crossmatch between our catalog and Hunt & Reffert (2023) varies from 15% (for NGC 6656; very close to the Galactic plane) to 95% based on the distance of the clusters from the Galactic plane, with a median crossmatch accuracy of 84%. NGC 104 is not studied by Hunt & Reffert (2023), hence we are unable to make any crossmatch with this cluster. The membership method by Hunt & Reffert (2023) is found to have a bias on whether the cluster located in a crowded or sparse field, which is not found in the methods of Vasiliev & Baumgardt (2021). In the top panel of Figure 12, we show CMDs of two GCs (NGC 3201 and NGC 288) to demonstrate the comparison of the cluster membership obtained from our method (blue), Vasiliev & Baumgardt (2021) in orange and Hunt & Reffert (2023) in green. The figure shows a good agreement between all three methods.
We find a better crossmatch with Hunt & Reffert (2023) for OCs. We find a crossmatch of more than 50% for all cluster (except UBC252 with 19%) and with median crossmatch value of 94%. The unmatched sources are mainly fainter than 18 mags in G band and more importantly, all BSSs in our sample are also classified as cluster members in Hunt & Reffert (2023). Dias et al. (2021) used the classic maximum likelihood approach (Dias et al. 2014; Monteiro et al. 2020) on Gaia-DR2 data to study 2000 open clusters for stars brighter than 18 mags in the G band. The comparison shows that the crossmatch fraction varies from 60% to 96% with a median value of 88%. The difference in the crossmatch is mainly found towards the fainter G magnitudes. When cross-matched only with our BSS sample, we find more than 97% are also classified as members in Dias et al. (2021). Cantat-Gaudin et al. (2020a) applied an artificial neural network in Gaia DR2 photometry down to G = 18 mags to estimate the cluster membership of 2000 OCs. Comparing our cluster members with stars brighter than 18 mags in G band, we find that crossmatch fraction varies from cluster to cluster with a range from 40% to 95%. There are mainly two reasons behind the differences in crossmatch fraction: Cantat-Gaudin et al. (2020a) used Gaia DR2 for the membership instead of DR3 with better astrometry and was unable to recover the cluster member properly towards the fainter magnitude limit. For example, there are no cluster member present with G17 mags for the cluster UBC 252, which leads to only 41% of our cluster member. While we consider the match for BSSs, we notice that out of 434 BSSs candidate in our catalog, 409 (94%) are also part of the cluster member in Cantat-Gaudin et al. (2020a). In the bottom panel of Figure 12, we present CMDs of two OCs (NGC 188 and Berkeley 39) as examples to display the comparison between our cluster membership (open blue box) to those obtained by Hunt & Reffert (2023) marked in orange points, Cantat-Gaudin et al. (2020a) marked in green cross, and Dias et al. (2021) in red points. The cluster memberships are found to be well-matched among the different methods. Jadhav & Subramaniam (2021) also provides an extensive catalog of BSSs candidates in OCs. However, they used the primary catalog of OCs by Cantat-Gaudin et al. (2020a). So, we did not perform any details crossmatch with Jadhav & Subramaniam (2021). Therefore, we find that our cluster membership matches well with the literature and almost all BSSs are also classified as members in the literature, which also validates the membership determination algorithm used in our study.
Appendix B Analysis
To obtain astrophysical parameters for BSSs we use three different approaches: i) Color–temperature relations; ii) Model isochrone fitting; iii) Spectroscopic data from Gaia DR3 (if available).
B.1 Color–temperature relations
Based on the infrared flux method (IRFM) from González Hernández & Bonifacio (2009), the effective temperature of a star can be estimated using the following relation.
where , is the dereddened color and is the metallicity. We use the coefficients from Mucciarelli et al. (2021), which are based on a Gaia EDR3 analysis of dwarf and giant stars. are taken from Carretta et al. (2009) and Dias et al. (2021) for GCs and OCs, respectively. For Pismis 3 and Trumpler 23, metallicities were taken from Bisht et al. (2022) and Overbeek et al. (2017), respectively.
B.2 Isochrone-fitting models
Previous studies have attempted to extract BSS parameters based on isochrone models such as JS21 and D23. In the first study the authors assume the mass of the BSSs to be approximately, the mass of a point at the ZAMS with the same magnitude. In the second study, they fit multiple isochrones with different ages and compare BSS positions with respect to isochrones. We take an approach similar to the second one. For GCs (OCs), we use multiple PARSEC isochrones (Bressan et al. 2012) from 200 Myr (100 Myr) in steps of 50 Myr (25 Myr) up to the cluster age found with the best-isochrone fitting, fixing the metallicity from the literature for each cluster as described earlier. We select the three closest points on the best-fit isochrone for each cluster and compute the average and rms for mass, luminosity, age, surface gravity and effective temperature for each BSS.
B.3 Spectroscopy from Gaia DR3
Gaia DR3 includes optical low–resolution spectra for over 220 million stars. As a brief summary, these spectra are measured with two instruments. One covering the blue–optical range (380–680 nm, measured with the “Blue Photometer” instrument) called BP and another covering infrared range (640–1050 nm, measured with the “Red Photometer”) called RP (De Angeli et al. 2023; Montegriffo et al. 2023). These BP/RP spectra, hereafter referred to as XP spectra, have a spectral resolution with parameters inferred from “classical” models (Andrae et al. 2023, and references therein), that are highly sensitive to errors in low-resolution spectra, which is the case for Gaia DR3 XP spectra. In another approach, Zhang et al. (2023), based on parameters obtained from higher-resolution spectra () from other surveys, cross-matched them with the Gaia DR3 spectra and created a machine-learning model that has the advantage of: i) obtaining astrophysical parameters from higher-resolution spectra, compare them with the XP spectra and adjust those parameters at lower spectral resolution; ii) being able to find various systematic errors that could not easily be noticed in the XP spectra alone. For this reason, we extract XP parameters (hereafter called XPP) from Zhang et al. (2023) for all our sample BSSs, if available, such as effective temperature and surface gravity. We cross-match the provided parameters with our selected BSSs using Gaia Source ID unique identifier. Finally, Zhang et al. (2023) provides a quality_flag parameter which, in short, is a number to quantify the reliability of the obtained parameters from spectra. The authors suggest a “basic reliability cut” to be at least (note, a lower number indicates more reliable parameters). We decided to only use stars with to ensure we extract only reliable results. More details for these parameters are provided in Section 4 of Zhang et al. (2023).
From the IRFM we find for GCs and for OCs. From isochrone fitting we find for GCs and for OCs. Only 537 out of 4399 BSSs ( of our sample) had XPP parameters measured for 279 stars in GCs and 258 stars in OCs. With this method, we find the mean BSS effective temperature of for GCs and for OCs. Figure 13 shows the distribution for , , and of all our sample BSSs. We have also estimated the parameters for YSSs and RSSs which will be available online as electronic tables. Their distributions are shown in Figure 15 and Figure 16 respectively.
To check the consistency of our results, we compare the values obtained from the three methods in Figure 14 against each other. The top figure is divided into two sub-panels, both of which are comparing derived with the isochrone fitting method against the values found with the IRFM (top–upper panel) and XPP (top–lower panel). The lower panel compares the values of latter two methods against each other. We only consider BSSs with an error in of less than in both components. The 1–to–1 relation is shown as a gray dashed line and compared with the data. We observe very good agreement between the methods, albeit with small level of systematics. We fit a linear model using LMFIT Python package777https://lmfit.github.io/lmfit-py/ to these relations considering full error propagation in all values, which are shown as a dot-dashed green lines. We also estimated pearson r-coefficients for all these relations in a monte carlo approach, and the estimated values are 0.93 (top panel), 0.76 (middle panel) and 0.74 (bottom panel). The errors correspond to 16th and 84th percentile values with respect to the median value. The r-coefficients suggest that estimated from different methods are strongly correlated. We obtain the following linear relations in Kelvin units:
From these results, we note that extracted from isochrone fitting and those derived from the color–temperature relations (IRFM) appear most consistent with each other up to , with little systematic offset towards higher IRFM temperature. A similar behaviour is observed for the XPP vs. isochrone method, where values are consistent within the scatter up to . Above that temperature, a small systematic offset towards lower XPP temperatures is observed. The bottom panel of Figure 14 shows the XPP vs. IRFM method comparison, which shows the largest scatter yet fairly consistent values for most BSSs, while the linear fit is pulled from 1–to–1 relation by some OC BSSs with temperatures above .
From these comparisons, we estimate that the systematic uncertainty between various methods is of the order K. The IRFM and isochrone methods consider the metallicity of the cluster, while the XPP method is inferior for the reasons described in Section B.3, which is reflected in the lower scatter and smaller systematic offset in the IRFM vs. isochrone method comparison (upper panel in Fig. 14). In addition, XPP parameters are not available for stars at the MSTO and we, therefore, are not able to extract using XPP. We, therefore, consider the IRFM and isochrone method from now as the bona-fide best estimates, keeping in mind the systematics described above.
Appendix C Astrophysical parameters for YSSs and RSSs
Similar to Figure 13, we display the extracted astrophysical parameters for stars classified as Yellow Straggler Stars (YSSs) and Red Straggler Stars (RSSs) in Figures 15 and 16, respectively.